Suppose that the information below describes the economy of Wonderland.

C = 300+0.8Yd
I = 300
G = 250
X = 300
M = 150 +0.6Yd
T = 0.25Yd

1) Suppose the investment (I) decreases with 50, calculate the new level of equilibrium income (Ye) by means of the MULTIPLIER
2) Calculate the tax income that the government of Wonderland receives at the equilibrium income level of question 1
1
Expert's answer
2020-11-20T09:41:39-0500

1) Y =C + I + G + X-M

Y = 300+0.8Yd+300+250+300-(150+0.6Yd)

300+300+250+300-150+0.8Yd-0.6Yd

Y = 1000+0.2Yd

New level of equilibrium when (I)= 300-50= 250

Y = C+I+G+X-M

300+0.8Yd+250+250+300-(150+0.6Yd)

300+250+250+300-150+0.8Yd-0.6Yd

950+0.2Yd


2) To get tax income

1000+0.2Yd = 950+0.2Yd

50 = 0.2Yd-0.2Yd

= 50

T = 0.25Yd

= 0.25×50

=12.5.

Please use the second answer.